OpenStudy (anonymous):

integral x tan^2 x dx solve using parts .. please help

OpenStudy (anonymous):

Let u = x dv = sec^2x dx du = dx v = tan x so by "integration by parts", you get x tan(x) - Integral of (tan x) dx = x (tan x) + ln/cos x/ + C Keep in mind that the integral of tan x is -ln/cos x/

OpenStudy (anonymous):

i thought the integral of tan x was ln/secx/+c

OpenStudy (anonymous):

ok, that will also work..either one is correct.

OpenStudy (anonymous):

Either one is correct; as they will both be correct after checking with differentiation.

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

so you will have x tan(x) - ln/sec x/ + C @karid

OpenStudy (anonymous):

but wouldn't dv = tan^2 x

OpenStudy (anonymous):

@karid BTW, if you ever forgot the integral of tan x...just write tan x as sin x/cos x...let u = cos x, du = -sinx dx so you get the integral of tan x dx = -ln/cos x/ + C

OpenStudy (anonymous):

@karid You are absolutely right...I thought it said sec^2(x) so tan^2(x) = sec^2(x) - 1 so integrate that...tan x - x sorry for the oversight.

OpenStudy (anonymous):

its ok ....thnx that makes more sense ..

OpenStudy (anonymous):

From Pythagorean Identity, 1 + tan^2(x) = sec^(2)x, so tan^2(x) = sec^2(x) - 1 @karid

OpenStudy (anonymous):

i know how to do the rest i forgot that i could change tan^2 x to sec^2 - 1 .. :)

OpenStudy (anonymous):

OK. All is cool.